OpenStudy (anonymous):

Torque Dynamics questions: 1. A large platter of food with total mass 3kg is dropped on to the middle of a spinning lazy-susan with spinning mass 1kg. Which has the greater change in angular momentum? A) the lazy-susan (less mass) B) the platter (more mass) C) their change in angular momenta are equal in magnitude What would change if the platter was dropped closer to the edge of the lazy-susan? A) the lazy-susan (less mass) B) the platter (more mass) C) their change in angular momenta are equal in magnitude

OpenStudy (anonymous):

2. A uniform beam of mass M and length L is attached to the wall by a hinge at one end and held horizontally. What is the linear acceleration of the tip of the beam if it is released? A) 0m/s2 B) g/2 C) g D) 3g/2 3. Two masses m1 > m2 are connected by string and hang on either side of a frictionless disk pulley with mass M and radius R. Where is the string tension greater? A) above m1 B) above m2 C) same everywhere 4. A yo-yo has inner radius R1, outer radius R2, and mass m distributed (approximately) uniformly, like a solid disk (ignoring the groove). What is the yo-yo's linear acceleration if released from rest and allowed to “roll” down the string? A) a = ( 1 + R2 / R1 ) g B) a = ( 1 + R1^2/R2^2)g C) a = g / ( 1 + R2^2/R1^2) D) a = g / ( 1 + R1 / R2 )

OpenStudy (anonymous):

1(a)

OpenStudy (anonymous):

1second part (b); 3 (A); in 2 the acceleration shoud be changing with time,not sure.. in 4 dunno how yo-yo roll..